A storage tank is 4.5 metre wide. It contains a liquid with density 750 kg/m^ 3 to a depth of 7 metre. Calculate the hydrostatic thrust on the side of the tank, and its position from the surface of the liquid.

1
Expert's answer
2021-11-08T08:50:14-0500

Area=4.5*7=31.5m2

Hydrostatic thrust, "F_c=\rho gh_cA"

"F_c=750*9.81*\frac{7}{2}*31.5"

"=811164.375"

"=811.16kN"

Location of force ,"y_p=\frac{2h}{3}"

"=\frac{2*7}{3}"

=4.667m